OpenStudy (anonymous):

I need help with this please x^3−3x^2+2x−17 on the interval [0,2] Find the values c that satisfy the conclusion of Rolle's theorem.

OpenStudy (anonymous):

good Question. :)

myininaya (myininaya):

Find f' Then solve f'(c)=0 for c

myininaya (myininaya):

\[f(x)=x^3-3x^2+2x-17\] What is f'(x)?

OpenStudy (anonymous):

3x^2 - 6x + 2

myininaya (myininaya):

Now \[f'(c)=3c^2-6c+2\] Set f'(c)=0 and solve for c \[3c^2-6c+2=0\]

OpenStudy (anonymous):

ok i got it thanks

myininaya (myininaya):

:)